Our company has 100,000 shares of $4 par 5% cumulative stock outstanding.

This year the board of directors declared $80,000 in dividends (the total

dividend, to be divided between preferred and common stock). Last year,

$6,000 in dividends were declared.

(1) "Dividends in arrears" is how much?

  

(2) How much TOTAL dividend does preferred stock receive this year?

(3) How much TOTAL dividend does common stock receive this year?
Answer #1

Annual Preferred Dividend = 100,000*4*5%= 20,000

Answers :

1) Dividend in Arrears (20000-6000) = 14,000

2) How much Dividend Paid to Preferred Stock (20,000+14000) = 34,000

3)How Many dividends Paid to Common Stock (80000-34,000) = 46,000